# Enigmatic factor of 4/3 in electromagnetic momentum of a moving spherical capacitor

### v2

@@ II Energy-momentum in the electromagnetic fields of a moving charged spherical capacitor
The momentum is zero in rest frame ${\cal K}^{\prime}$ as the magnetic field is throughout zero in that frame.
@@
$\frac{{\rm d}{U}}{{\rm d}t}$ $=$ $\frac{{\rm d}{\bf P}}{{\rm d}t}\cdot{\mbox{\boldmath$\beta$}}{c}\>,$ (9)
for the same energy ${U}^{\prime}$ in the rest frame. Here we have used for a one-dimensional motion, ${{\rm d}{\gamma}}/{{\rm d}t}=\gamma^{3}\mbox{\boldmath$\beta$}\cdot{\rm d}\mbox{\boldmath$\beta$}/{\rm d}t$ and ${{\rm d}({\gamma}\mbox{\boldmath$\beta$}})/{{\rm d}t}=\gamma^{3}{\rm d}\mbox{\boldmath$\beta$}/{\rm d}t$.
@@
$=$ $\gamma\left(1+\frac{\beta^{2}}{3}\right)\frac{Q^{2}}{2C}=\frac{4}{3}\,{U}^{\prime}\gamma-\frac{{U}^{\prime}}{3\gamma}\>.$
${\bf P}_{\rm f}$ $=$ $\frac{4}{3}\frac{\gamma\mbox{\boldmath$\beta$}}{c}\frac{Q^{2}}{2}\left(\frac{1}{a}-\frac{1}{b}\right)=\frac{4}{3}\frac{\gamma\mbox{\boldmath$\beta$}}{c}\frac{Q^{2}}{2C}$ (13)
$=$ ${U}_{\rm f}\frac{\mbox{\boldmath$\beta$}}{c}+\frac{{U}^{\prime}}{3\gamma}\frac{\mbox{\boldmath$\beta$}}{c}\>.$
There are a couple of noteworthy points here. First, compared to Eq. (8), field momentum has an additional term, the second term on the right hand side, in Eq. (13). Also from Eqs. (12) and (13), we get
$\frac{{\rm d}{U}_{\rm f}}{{\rm d}t}=\frac{{\rm d}{\bf P}_{\rm f}}{{\rm d}t}\cdot{\mbox{\boldmath$\beta$}}{c}-\,\frac{{U}^{\prime}}{3}\frac{{\rm d}}{{\rm d}t}\left(\frac{1}{\gamma}\right)\,,$ (14)
which, as compared to Eq. (9), has an additional term, the second term on the right hand side.
@@ III.2 Energy contribution from electromagnetic forces during work done against Lorentz contraction
Suppose in the lab frame ${\cal K}$, where the capacitor system is moving with velocity $\beta=v/c$ along $x$-axis, the capacitor gets a velocity increment $\Delta\beta$ along $x$-axis, and a corresponding increment in the Lorentz factor $\Delta\gamma\approx\gamma^{3}\beta\Delta\beta$, in a small time interval $\Delta t$. Because of the increase in the Lorentz factor, the distance between $R_{1}$ and $R_{2}$ decreases in frame ${\cal K}$ by

### v3

@@ II Energy-momentum in the electromagnetic fields of a moving charged spherical capacitor
II.1 Energy-momentum computed from a 4-vector Lorentz transformation from the rest frame
@@
The electromagnetic momentum $\bf P^{\prime}$ is zero in rest frame ${\cal K}^{\prime}$ as the magnetic field is zero throughout in that frame. If the system has an acceleration ${\rm d}\mbox{\boldmath$\beta$'}/{\rm d}t^{\prime}$ in the instantaneous rest frame, then we can write ${\rm d}{\bf P}^{\prime}/{\rm d}t^{\prime}=(U^{\prime}/c){\rm d}\mbox{\boldmath$\beta$'}/{\rm d}t^{\prime}$. The rate of change of energy ${U}^{\prime}$ in the rest frame is nil, i.e., ${\rm d}U^{\prime}/{\rm d}t^{\prime}=0$.
@@
Because of an acceleration ${\rm d}\mbox{\boldmath$\beta$}/{\rm d}t$ of the system in the lab frame, we have
$\frac{{{\rm d}{\gamma}}}{{\rm d}t}$ $=$ $\gamma^{3}\mbox{\boldmath$\beta$}\cdot\frac{{\rm d}\mbox{\boldmath$\beta$}}{{\rm d}t}\,,$ (9)
$\frac{{{\rm d}({\gamma}\mbox{\boldmath$\beta$}})}{{\rm d}t}$ $=$ $\gamma^{3}\Big{(}\mbox{\boldmath$\beta$}\cdot\frac{{\rm d}\mbox{\boldmath$\beta$}}{{\rm d}t}\Big{)}\mbox{\boldmath$\beta$}+\gamma\frac{{\rm d}\mbox{\boldmath$\beta$}}{{\rm d}t}\,.$ (10)
From Eq. (6) and Eq. (7), with the help of Eqs. (9) and (10), we can write
$\frac{{\rm d}{U}}{{\rm d}t}$ $=$ $U^{\prime}\frac{{\rm d}{\gamma}}{{\rm d}t}=U^{\prime}\gamma^{3}\mbox{\boldmath$\beta$}\cdot\frac{{\rm d}\mbox{\boldmath$\beta$}}{{\rm d}t},$ (11)
$\frac{{\rm d}{\bf P}}{{\rm d}t}$ $=$ $U^{\prime}\frac{{\rm d}({\gamma}\mbox{\boldmath$\beta$})}{c\>{\rm d}t}=U^{\prime}\gamma^{3}\Big{(}\mbox{\boldmath$\beta$}\cdot\frac{{\rm d}\mbox{\boldmath$\beta$}}{c\>{\rm d}t}\Big{)}\mbox{\boldmath$\beta$}+U^{\prime}\gamma\frac{{\rm d}\mbox{\boldmath$\beta$}}{c\>{\rm d}t}.$ (12)
Then from Eq. (11) and Eq. (12), using $\gamma^{2}\beta^{2}+1=\gamma^{2}$, we get
$\frac{{\rm d}{U}}{{\rm d}t}=\frac{{\rm d}{\bf P}}{{\rm d}t}\cdot{\mbox{\boldmath$\beta$}}{c}\>.$ (13)
This can be understood in another way. Because of the force $\bf F={{\rm d}{\bf P}}/{{\rm d}t}$ in the lab frame ${\cal K}$, where the system is moving with a velocity ${\bf v}=\mbox{\boldmath$\beta$}{c}$, energy of the system is changing at a rate ${\bf F\cdot v}$.
II.2 Energy-momentum computed directly from electromagnetic fields of a moving charged capacitor
@@
$=$ $\gamma\left(1+\frac{\beta^{2}}{3}\right)\frac{Q^{2}}{2C}=\frac{4}{3}\,{U}^{\prime}\gamma-\frac{{U}^{\prime}}{3\gamma}\>,$
${\bf P}_{\rm f}$ $=$ $\frac{4}{3}\frac{\gamma\mbox{\boldmath$\beta$}}{c}\frac{Q^{2}}{2}\left(\frac{1}{a}-\frac{1}{b}\right)=\frac{4}{3}U^{\prime}\frac{\gamma\mbox{\boldmath$\beta$}}{c}\>.$ (17)
From Eq. (16) and Eq. (17), with the help of Eqs. (9) and (10), we can write
$\frac{{\rm d}{U_{\rm f}}}{{\rm d}t}$ $=$ $\frac{4}{3}U^{\prime}\frac{{\rm d}{\gamma}}{{\rm d}t}-\,\frac{{U}^{\prime}}{3}\frac{{\rm d}}{{\rm d}t}\left(\frac{1}{\gamma}\right)$ (18)
$=$ $\frac{4}{3}U^{\prime}\gamma^{3}\mbox{\boldmath$\beta$}\cdot\frac{{\rm d}\mbox{\boldmath$\beta$}}{{\rm d}t}-\,\frac{{U}^{\prime}}{3}\frac{{\rm d}}{{\rm d}t}\left(\frac{1}{\gamma}\right)\>,$
$\frac{{\rm d}{\bf P}_{\rm f}}{{\rm d}t}$ $=$ $\frac{4}{3}U^{\prime}\frac{{\rm d}({\gamma}\mbox{\boldmath$\beta$})}{c\>{\rm d}t}$ (19)
$=$ $\frac{4}{3}U^{\prime}\gamma^{3}\Big{(}\mbox{\boldmath$\beta$}\cdot\frac{{\rm d}\mbox{\boldmath$\beta$}}{c\>{\rm d}t}\Big{)}\mbox{\boldmath$\beta$}+\frac{4}{3}U^{\prime}\gamma\frac{{\rm d}\mbox{\boldmath$\beta$}}{c\>{\rm d}t}\>.$
Then from Eq. (18) and Eq. (19), using $\gamma^{2}\beta^{2}+1=\gamma^{2}$, we get
$\frac{{\rm d}{U}_{\rm f}}{{\rm d}t}=\frac{{\rm d}{\bf P}_{\rm f}}{{\rm d}t}\cdot{\mbox{\boldmath$\beta$}}{c}-\,\frac{{U}^{\prime}}{3}\frac{{\rm d}}{{\rm d}t}\left(\frac{1}{\gamma}\right)\,.$ (20)
There are a couple of noteworthy points here. Compared to Eq. (13), Eq. (20) has an additional term, the second term on the right hand side. As we will demonstrate in section (III.2, Eq. (31)), this additional term in Eq. (20) represents the rate of work being done during Lorentz contraction of the system against the electromagnetic forces in the charged system. Also, using Eq. (16) to write Eq. (17) as
${\bf P}_{\rm f}$ $=$ ${U}_{\rm f}\frac{\mbox{\boldmath$\beta$}}{c}+\frac{{U}^{\prime}}{3\gamma}\frac{\mbox{\boldmath$\beta$}}{c}\,.$ (21)
we note that compared to Eq. (8), there is an additional term in field momentum, the second term on the right hand side, in Eq. (21). It is this term, whose origin will be explored in detail in section (III.1, Eq. (25)), that is responsible for the enigmatic factor of 4/3 in electromagnetic momentum of a moving charged system.
@@ III.2 Energy contribution from electromagnetic forces during work done against Lorentz contraction
Suppose in the lab frame ${\cal K}$, where the capacitor system is moving with velocity $\beta=v/c$ along $x$-axis, the capacitor gets a velocity increment $\Delta\beta$ due to an acceleration component along the $x$-axis, and a corresponding increment in the Lorentz factor $\Delta\gamma\approx\gamma^{3}\beta\Delta\beta$, in a small time interval $\Delta t$. Because of the increase in the Lorentz factor, the distance between $R_{1}$ and $R_{2}$ decreases in frame ${\cal K}$ by